Anjali wants to make 10.5 qt. of 50% sugar solution by mixing together a 70% sugar solution and a 10% sugar solution. How much of each solution must she use?
Answer by mananth(16946)

Anjali wants to make 10.5 qt. of 50% sugar solution by mixing together a 70% sugar solution and a 10% sugar solution. How much of each solution must she use?
Let 10% soln be x
70*sol will be 10.5 -x
10%---------------------70%------------------50%
x-----------------------10.5-x ---------------10.5 qt.
0.1x+ 0.7(10.5-x)=0.5*10.5
0.1x+7.35-0.7x=5.25
-0.6x=-2.1
x=3.5 qt of 10% solution
10.5 - 3.5 = 7 qt of 70% solution